# Assembly Theory Reduced to Shannon entropy and Rendered Redundant by Naive Statistical Algorithms

### v7

@@ 3 Discussion
Our results have proven that an object with a low assembly index has high $\mathbf{LZ}$ compressibility, and therefore low Entropy. In the opposite direction, an object with a high assembly index will have low $\mathbf{LZ}$ compressibility, and therefore high Entropy. For every assembly step, there is a block repetition. Therefore, counting the number of assembly steps is equivalent to counting the number of repetitions of each word in a dictionary.
@@
AT and its assembly index can, therefore, only reproduce the capabilities already shown and known for Entropy-based measures and $\mathbf{LZ}$ algorithms, both of which are approximations to algorithmic complexity.
@@
To the extent that AT may have a pedagogical value in its peculiar application to molecular complexity, this has unfortunately remained unclear due to the obfuscatory effect of the hyperbole surrounding the introduction of AT by its authors, amplified by the scientific media, with AT being touted as a theory of everything able to unify physics and biology, explain and characterise life on Earth and beyond, explain matter, time, revolutionise medicine and drug discovery, and even explain the expansion of the universe and the phenomenon of cosmic inflation [13, 14, 6], which is an open problem in cosmology.
@@
It may appear to be an advantage of AT that it may represent a transparent-box algorithm. This is in fact not the case, since a different heuristic of it has been introduced in every new paper. Most of the time, this forces us to take their empirical results at face value, thus giving them the greatest advantage while hindering reproducibility and refutation processes.
However, we still match or outperform their results with naive statistical algorithms. We have formally proved that any implementation of the copy-number argument, hence any measure based on the assembly index (Ai) idea, is equivalent to Shannon entropy rate and dictionary-based statistical compression algorithms such as those in the $\mathbf{LZ}$ family underlying PNG and ZIP [4].
In fact, the underlying idea of $\mathbf{LZ}$ algorithms, such as LZW or the assembly index calculation method, is to find redundancy in the data, as Shannon entropy does. They achieve this by counting exact copies of smaller objects in a larger object. The fact that these decades-old functions are being renamed does not add new capabilities to the wide range of applications of compression algorithms and Shannon entropy, many of which have already been reported before [12, 22, 34] without the hyperbolic claims.
@@
We have been very critical of the use and abuse of compression as a complexity measure and have been addressing this issue for more than a decade. When it comes to Assembly Theory, we think that the limitations of statistical compression cannot be rectified with another (disguised) compression algorithm of exactly the same statistical nature. Therefore, we are not advocating for statistical compression [32]. In fact, for more than a decade we have worked to move the field away from statistical compression and simplistic measures of complexity like AT.
@@
The discussion around AT has been plagued by hype and obfuscation. However, we recognise the effort on the part of the authors to offer a more formal framework for discussion of the details relating to their index in [1], unlike previous attempts with even more serious mistakes [23], ranging from an unnecessary proof of computability, (as if $\mathbf{LZ}$ or Shannon entropy needed proofs of computability) to baseless claims that it was a generalisation of algorithmic (Solomonoff-Kolmogorov-Chaitin) complexity [20].
The authors have continued to make increasingly hyperbolic claims (including the most recent claims about AT’s ability to explain the expansion of the universe [6]), to fail to cite and benchmark against previous work published in the literature, and to disseminate misinformation about theoretical computer science and information theory to the general public, even advocating that such contradictions or deleterious practices might be useful as motivation for future work [25].
@@ 3.1 Unfounded new arguments based on stochasticity
Despite overwhelming evidence that Assembly Theory (AT) does not offer any fundamental advantage over Shannon entropy and other methods, its proponents have never tested it against any other index or measure and continue to make unfounded, false, and uninformed claims about both their own and other measures.
The authors of Assembly Theory continue to act surprised and reject the suggestion that Assembly Theory and their assembly index are, willy nilly, data compression schemes, which is what we have shown they are [4, 2] regardless of the authors’ personal opinions.
One of their objections is that compression is a modern concept from computer science, hence anachronistic vis-à-vis the prebiotic phenomena they wish to quantify, computers not having predated humans. This argument makes no sense from any perspective. As scientists we are meant to analyse nature with the best tools humans have created, whether drawn from mathematics or computer science. Moreover, their core concepts of copy, number, or graph do not predate humans or biotic systems on Earth either. The authors of AT tend to confound the way an algorithm may be described with how nature may instantiate it [4]. Assembly Theory’s performance –including measuring or estimating the assembly index and assembly number, or modelling assembly spaces– can be replicated by many types of mathematical formalisms, and therefore by things that appeared much later than the advent of life on Earth.
For example, as we have formally shown in [4], exploring all the possible assembly spaces is equivalent to exploring a subspace of context-free grammars (CFG). The assembly index is equivalent to the size of such compressing CFGs.
This is because for every assembly (sub)space there is a CFG that not only generates the same objects but also is isomorphic vis-a-vis the particular multigraph put forward by Assembly Theory. In addition, the minimum number of required assembling steps (or “assembly time”) is exactly the same as the height of such a CFG tree.
In fact, because all of the methods of Assembly Theory are computable (even if the calculation of the assembly index may be intractable), there is an unrestricted grammar that generates–via a computable process– all the possible assembly spaces. This in turn further uncovers misleading arguments [16, 1] involving the requirement of Turing machines, determinism vs indeterminism, computability vs stochasticity, thus involving equivalence classes of mathematical functions closed under reductions [4]. In contrast to being defined for a particular type of computable process only, e.g. deterministic Turing machines, such arguments overlook equivalence classes for which algorithmic complexity is known to be invariant. In order to calculate the assembly index and the assembly number, instead of the implemented algorithms or mathematical functions used in AT, one can for example (whether abstractly or physically) implement a myriad distinct formal processes, such as lambda calculus, elementary cellular automata (ECA), Conway’s game of life, any primitive recursive function together with the (unbounded) $\mu$-operator, unrestricted grammars, nondeterministic Turing machines (NTM), or even some (proper) classes of probabilistic Turing machines (PTM) that are stochastic by design (or assumption).
Shannon entropy was, however, intrinsically designed to deal with uncertainty. It is the quintessential function of uncertainty as a probability distribution multiplied by the past occurrence of the random variable. It counts the number of elements repeated in an object to quantify the statistical likelihood of the next element. Everything that AT does with its assembly index, Shannon entropy can do, and everything that Shannon entropy cannot do, Assembly Theory cannot, because they are fully equivalent [4].
@@ 3.2 Unfounded new arguments based on Time complexity
Another key flaw in the latest arguments from AT’s authors [1] lies in their reliance on inconsistent heuristics, which change from paper to paper. These heuristic tweaks, designed to maintain marginal differences from Shannon entropy or LZ compression, amount to shifting goalposts. Despite theoretical and empirical proofs demonstrating that the assembly index is fundamentally equivalent to Shannon entropy, the authors persist in claiming novelty by pointing to small scatter deviations in data. These deviations, however, arise not from theoretical innovation but from arbitrary algorithmic inefficiencies.
Despite the authors’ more modest reformulation of previously established findings in the article of interest [1], the AT story has, surprisingly, remained in the news cycle for the last few years, fuelled by the authors’ unfounded claims (e.g. [13, 14, 6]) and media presence.
We recognise that they have backtracked on some of their false claims, now admitting that it was always incorrect to assert that Assembly Theory is not a generalisation of algorithmic complexity [1], as we proved in [4]. They have also retracted their claim that the implementation of their assembly index does not require a Turing machine—-when, in fact, it does, because it remains a computer program that must be executed on a digital computer.
However, one of the most baffling arguments most recently presented by the authors of AT [1], is their insistence that their algorithm is different because it belongs to a different Time complexity class.
@@
While time complexity is an important distinguishing feature of an algorithm in computer science, in science, and for scientific purposes, the strength of a method lies in its explanatory power and predictive accuracy–not in its inefficiency or efficacy. Yet, the argument of AT is that they are worse and slower and therefore are different. While this argument borders on the ludicrous, it is also wrong. Indeed, slow implementations of $\mathbb{LZ}$ algorithms exist that converge faster to Entropy rate. Even compressing tools like gzip have compression levels as a parameter ranging from 1 to 9, with 1 being the fastest compression (but with a lower compression ratio) and 9 being the slowest compression (but with the highest compression ratio); the default level is usually set to 6, striking a balance between speed and compression quality, but at 9 gzip is very inefficient and eventually intractable for very long sequences and large objects. Even though at level 1 it still guarantees convergence to Shannon Entropy rate in the limit of an (infinite) object–the same eventual convergence for level 9–the rate of convergence varies drastically. In other words, one can implement a Shannon entropy routine that can be as fast or slow as needed as a function of sliding window size and total object size.
We have proven that applying AT and their assembly index is simply applying Shannon entropy with extra steps. One can reproduce all their results by directly applying Shannon entropy for all their experiments. Something the authors have always avoided and rejected to do as a comparison and basic control experiment. AT offers no additional explanatory power beyond what Shannon entropy-based explanations can already provide.
The authors go to great lengths to provide proofs of the NP-completeness of their assembly index (AI or Ai), in a futile exercise of fake rigour, as if computational complexity justified their theory’s novelty, explanatory or predictive power. These authors’ arguments amount to what we have identified as the we-are-different-because-we-are-worse fallacy.
3.3 False arguments from misleading experiments
@@
We have here demonstrated that to separate molecules produced by living and non-living systems, Marshall et al. [5] used a linear model to estimate MA (Ai) values from MS2 peak data. Then, we showed that this linear model (which is the core of the methodology in [5]) is easily replaced by any other model that correlates any metric with the MS2 peak data.
@@
The authors of Assembly Theory are mistaken in claiming to have been the first to report on the separation of organic from inorganic compounds using complexity measures. We reported the separation of organic from inorganic compounds in 2018 [12] (on a complete database with more than 15 000 molecules) while they did so in 2021 [5] (on a small set of cherry-picked molecules). In 2018, we also reported connections to selection and evolution [22], which they did in 2023 [3]. All this without the hyperbolic claims that have unjustifiably followed Assembly Theory.
@@
We have here shown that it is possible to build a linear model, with similar predictive performance, to estimate the InChI string lengths of the molecules from MS2 peak data. In the end, following the same rationale as the authors of AT, we could claim that LZW (applied on InChI strings) and even the length of InChI strings could be measured experimentally and could also be used to separate molecules produced by living and non-living systems. As a matter of fact, their rationale leads us to say that non-living systems do not produce molecules whose estimated InChI string lengths are above 100 characters. This is trivially false.
@@
While some complexity measures may be similar or may overlap in scope or method, as stated in [1], one has to demonstrate any added value or justifiable advantage in predictive or explanatory power of the explicandum. Assembly Theory is neither predictive nor explanatory. The results shown so far are easily reproducible or outperformed, and the alleged explanatory power is based on very empirical results that do not indicate that their peculiar assembly processes are in any way favoured. The connections to selection or evolution are only speculative and are based on circular arguments [4].
@@
The authors of AT have also failed to address the mathematical proof we offered, which shows that their assembly index (Ai), assembly pathways, and all associated ideas based on counting identical copies in experimental or ‘physical’ data with a computable representation are mathematically equivalent to Shannon entropy [4].
Ultimately, AT’s inability to demonstrate superiority over existing measures, combined with its refusal to engage in rigorous comparisons, undermines its validity. Its claims to measure complexity in a novel way remain unsupported, and its insistence on dismissing well-established metrics such as Shannon entropy and LZ compression on the basis that AT explains selection and evolution–despite a lack of any evidence–cannot withstand scrutiny and must be ignored.
Assembly theory and its index offer no evidence to suggest that it is more related to evolution or selection than any other measure.
Indeed, they complain that Shannon entropy or LZ compression can only be applied to bits, but this is mistaken, too; it can be applied to any object and on any basic units, from vertices in a network to peaks from a mass spectral file.
@@
The authors of AT have suggested that the statistical algorithms we used in our previous articles [2, 4] successfully replicated all of their results because we did not control for molecular length. However, this was precisely our original criticism of their own methods. Our findings show that their assembly index is entirely dominated by molecular length, as we have demonstrated in this contribution, proving that their separation between organic and non-organic compounds was determined exclusively by length.
@@
It is illogical that the authors now appear to be reversing the argument, accusing us of the very oversight that they themselves have demonstrably failed to account for in the first place. Specifically, they failed to control for the most basic features and failed to compare their index with other measures of complexity. Had they done so, they would have immediately found that their whole groundbreaking theory was unable to produce any new results, that the results were entirely driven by Shannon entropy and compression techniques previously reported as being able to separate organic from nonorganic compounds.
@@
The authors of AT turned the tables again when we accused them of using a popular statistical compression disguised as an ‘assembly index’, then accusing us of suggesting that statistical data compression was equal to or better than Ai for characterising life. Instead, our argument is that none of these algorithms, including Shannon entropy, statistical compression, and Ai as a derived Shannon entropic index and compression algorithm, can characterise life because they are too trivial and ill-equipped to do so.

### v10

@@ 3 Discussion
Although Assembly Theory has been promoted in some outlets as a scientific breakthrough, driven in part by press releases, it has yet to demonstrate substantive contributions beyond established concepts in complexity and information theory. This paper, in continuing to refute hyperbolic claims made on behalf of AT, serves to correct several misconceptions about computability, algorithmic complexity, and compressibility.
@@
Our results confirm that an object with a low assembly index has high LZ compressibility and therefore low Shannon entropy, and an object with a high assembly index has low LZ compressibility and therefore high Shannon entropy. AT and its assembly index can therefore only reproduce the capabilities already known for entropy-based metrics, for which LZ algorithms are already optimal estimators.
@@
To the extent that AT may have a pedagogical value in its peculiar application to molecular complexity, this has unfortunately remained unclear due to the hyperbole surrounding the introduction of AT by its authors touted as a theory capable of even explaining the expansion of the universe and cosmic inflation [33, 34, 3].
@@
It may appear to be an advantage of AT that it appears as a transparent-box algorithm. This is in fact not the case, since a different heuristic of the same metric has been introduced in every new paper. Most of the time, this forces us to take their empirical results at face value, thus giving them the greatest advantage while hindering reproducibility. Moreover, one can always open the box of a statistical compressor and force the definition of dictionary elements to make the compressor fully transparent.
@@
We have been very critical of the use and abuse of compression as a complexity measure before, and we have been addressing this issue for more than a decade [31]. When it comes to Assembly Theory, we think that the limitations of statistical compression cannot be rectified with another compression algorithm of exactly the same limited nature.
In this paper, we are not advocating for statistical compression [31]. For more than a decade we have worked to move the field away from statistical compression and simplistic measures of complexity like AT towards recursive ones capable of more than statistics, like CTM and BDM.
@@
We recognise the effort of the authors of AT to offer a more formal framework for discussion of the details relating to their index in [4], unlike previous attempts [26] that lacked any attempt to compare and engage with current literature and methods, moving away from other previous mistakes made by the same senior authors, ranging from providing an unnecessary proof of computability [26] to baseless claims that AT and Ai were a generalisation of algorithmic complexity (Solomonoff-Kolmogorov-Chaitin) as voiced in [16]. However, the discussion around AT has historically been plagued by hype and obfuscation since its inception.
@@ 3.1 Unfounded new arguments based on stochasticity
One of the authors of AT’s objections is that compression is a modern concept from computer science, hence anachronistic vis-à-vis the prebiotic phenomena, computers not having predated humans, they say. This argument makes no sense from any perspective. As scientists we are meant to analyse nature with the best tools humans have at their disposal, whether drawn from mathematics or computer science. One does not need to justify mathematics as human-made to apply it as a universal tool. Moreover, the core concepts of AT such as ‘copy’, ‘number’, or ‘graph’ are also human constructions and do not predate biotic systems on Earth either. The authors of AT tend to confound the way an algorithm can be described with how nature may instantiate it [17]. Describing a hypothesis is not the same as explaining the phenomenon for which they have proposed a hypothesis, especially given Assembly Theory’s poor performance compared to ‘a-vis Shannon entropy or LZW offering no advantage.
For every assembly (sub)space there is a compressing grammar that not only generates the same objects but also is isomorphic vis-a-vis the particular multigraph put forward by Assembly Theory [17].
In fact, because all of the methods of Assembly Theory are computable (even if the calculation of the assembly index may be intractable), there is an unrestricted grammar that generates—via a computable process—all the possible assembly spaces.
Arguments around the unsuitability of Turing machines or algorithmic complexity to deal with uncertainty and stochasticity [15, 4] overlook the fact that algorithmic probability is the accepted mathematical Bayesian theory of optimal induction/abduction. We elaborate more on this in 3.5.
Furthermore, Shannon entropy is the quintessential metric to deal with stochasticity, which is equivalent to AT and its assembly index, therefore, making redundant the use of another metric [17] to deal with the same challenge.
@@ 3.2 Unfounded new arguments based on Time complexity
Another seminal flaw in the latest arguments from AT’s authors [4] lies in their reliance on inconsistent heuristics, which change from paper to paper. These heuristic tweaks, designed to maintain marginal differences from Shannon entropy or LZ compression, amount to shifting goalposts. Despite theoretical and empirical proofs demonstrating that the assembly index is fundamentally equivalent to Shannon entropy, the authors persist in claiming novelty by pointing to small scatter deviations in the data. However, these deviations arise not from theoretical innovation but from arbitrary algorithmic inefficiencies.
@@
Although time complexity is an important feature of an algorithm in computer science, in science, and for scientific purposes, the strength of a method lies in its explanatory power and predictive accuracy, not in its inefficiency or efficacy. However, the new argument of the AT group is that they are worse and slower and, therefore, are different. Although this argument borders on sounding absurd as a positive attribute, it is also wrong. There are already slow implementations of LZ algorithms that may converge faster or not to Shannon entropy rate. Even compression tools like gzip have compression levels as input parameter ranging from 1 to 9, with 1 being the fastest compression (but with a lower compression ratio) and 9 being the slowest compression (but with the highest compression ratio) depending on variations such as multiple passes through the object.
The AT authors’ new argument based on time complexity leads to a contradiction. They dismiss Kolmogorov complexity for its reliance on the Turing machine formalism which it does not, yet appeal to time complexity—a framework entirely dependent on that same formalism—as a new defending argument in favour of Ai.
Time complexity theory provides a powerful framework for classifying algorithms under the Turing machine model, but its relevance to natural processes is minimal. Complexity classes such as P, NP, and PSPACE are entirely model-dependent and assume the sequential execution of operations under the conventions of the Turing formalism. Even minor changes to the computational substrate can radically alter time-complexity results. For example, adopting a parallel architecture such as a cellular automaton can collapse exponential processes into linear or constant-time ones. Unlike algorithmic (Kolmogorov) complexity, which remains invariant up to an additive constant across all computational models, time complexity does not survive architectural changes.
A simple physical example illustrates this disconnect. Consider the task of sorting straws by length. In standard algorithmic analysis, comparison-based sorting is bounded below by $O(n\log n)$ operations. In the physical world, however, one may grasp a bundle of straws and tap them against a table, causing them to align by length in a single action—effectively an $O(1)$ sorting procedure. This highlights that time complexity classes are artefacts of a particular abstract model, not descriptions of physical resource use. Physical systems exploit parallelism, dynamics, and material constraints that are not captured by asymptotic complexity classes.
The contradiction in the defence of Assembly Theory (AT) and the assembly index (Ai) is therefore stark. AT mistakenly dismisses Kolmogorov complexity on the grounds that it relies on the Turing machine formalism, while simultaneously appealing to time complexity classes to distinguish Ai from compression. Yet, by virtue of requiring Turing universality, unlike time-complexity, Kolmogorov complexity is independent of the formalism or architecture of the universal constructor and its Invariance Theorem, while time complexity is entirely dependent upon it. Invoking time complexity as a defence of Ai while rejecting Kolmogorov complexity is thus not only misplaced but also reveals a repeated conflation and misunderstanding of the foundations of these areas within AT.
3.3 False arguments from incomplete experiments
@@
We have demonstrated that to separate molecules produced by living and nonliving systems, Marshall et al. [1] used a linear model to estimate MA (Ai) values from MS2 (mass spectral) peak data. Then, we showed that this linear model (which is the core of the methodology in [1]) is easily replaced by any other model that correlates any metric with the MS2 peak data.
@@
The authors of Assembly Theory are mistaken in claiming to have been the first to report on the separation of organic from inorganic compounds using complexity measures. We reported this separation in 2018 [22] while they did so in 2021 [1] (on a small set of molecules). In 2018, we also reported connections to selection and evolution [29], which they did in 2023 [2]. All this without the hyperbolic claims that have unjustifiably followed Assembly Theory.
@@
We have shown here that it is possible to build a linear model, with similar predictive performance, to estimate the InChI string lengths of the molecules from MS2 peak data.
@@
Although some complexity measures may be similar or may overlap in scope or method, as stated in [4], one has to demonstrate any added value or justifiable advantage in the predictive or explanatory power of the explanation. Assembly Theory is neither predictive nor explanatory. The results shown so far are easily reproducible or outperformed, and the alleged explanatory power is based on very empirical results that do not indicate that their peculiar assembly processes are in any way favoured. The connections to selection or evolution are only speculative and are based on circular arguments [17]. We further provide more examples in this work.
@@ 3.4 The Assembly index as a metric
The authors of AT have suggested that the statistical algorithms we used in previous articles [5, 17] that successfully reproduced all of their results did so because we did not control for molecular length. However, this was precisely our original criticism of their methods. Our findings show that their assembly index is entirely dominated by molecular length, as we have demonstrated in this contribution, proving that their separation between organic and nonorganic compounds was determined exclusively by length.
@@
Our argument has always been that the authors of AT did not control for the most basic features to compare their index with other measures of complexity. Had they done so, they would have immediately found that their whole groundbreaking theory was unable to produce any new results, that the results were entirely driven by Shannon entropy and compression techniques.
@@
The authors of AT also used the argument that we are suggesting that statistical data compression can characterise life. Our argument is that none of these algorithms can, including Shannon entropy and statistical compression of which Ai is a derived index and another statistical compression algorithm.
The authors suggest that Shannon entropy and compressibility are not relevant “metrics” in their domain. This is incorrect and reveals a lack of engagement with decades of empirical work using entropy and LZW in fields ranging from spam detection to genomics and evolutionary analysis.
3.5 The illusion of physical grounding
Unlike the claims made by the authors of AT in [4], Kolmogorov complexity does not require the construction of a universal Turing machine (UTM). The machine only contributes a fixed additive constant that is discounted in all cases. By contrast, the assembly index is also computed symbolically: it is implemented as code running on a digital computer, itself a universal machine, without anyone needing to build the computer from scratch. If their argument were applied consistently, they would face the same question they pose to Kolmogorov complexity: Who or what is assembling the blocks? The answer, whether “nature”, the “laws of physics or biology”, or “the computer”, is the same in all cases.
The attempt to dismiss Kolmogorov complexity on the grounds that it presupposes a universal Turing machine rests on a basic misunderstanding of the measure itself. In algorithmic information theory, the choice of universal Turing machine is irrelevant beyond a fixed additive constant. The UTM was introduced precisely to demonstrate that one need not account for the particular machine at all—the concept of universality guarantees that the substrate does not matter, so long as it is computationally complete. This is why it has long been possible to speak meaningfully about the Kolmogorov complexity of DNA sequences, molecular graphs, or any other object: the definition applies to the description, not to the physical act of building either the object or the machine.
The historical trajectory further undermines their claim. Turing’s universal machine was not conceived as an abstraction detached from reality but as a formalisation of the mundane practice of human “computers” calculating step by step with paper and pencil. The abstraction was then added to show that the details of the substrate were immaterial. To invert this order, demanding physical instantiation of the abstraction, is to miss both the logic and the foundations of computability theory.
Invoking von Neumann to demand prior physical construction also misconstrues von Neumann’s project. Von Neumann’s “universal constructor” was an existence proof carried out entirely in a formal medium. The aim was to establish the logical sufficiency of a description-directed constructor for self-reproduction and open-ended construction within an abstract substrate—not to require that such a device be physically built before one can reason about construction or complexity.
3.6 Physicality v Digital Representations
One of the most persistent rhetorical devices in this defence is the claim that AT operates on “physical data”, whereas compression algorithms are said to deal only with “abstractions”. This narrative has been used to elevate AT as fundamentally different from, and supposedly superior to, established information-theoretic methods. However, this distinction is both intellectually confused and technically incorrect.
The rhetoric of “physicality” obscures the fact that both their assembly index and Kolmogorov complexity operate in the same way: as symbolic measures, implemented in computer programs running on universal machines, applied to encoded descriptions of objects rather than to the physical objects themselves. Ultimately, both Kolmogorov complexity and the assembly index are operationalised as algorithms executed on universal machines acting upon symbolic encodings, and Assembly Theory may be understood as a restricted instance of Kolmogorov complexity, one that effectively corresponds to the programmatic implementation of Shannon entropy.
The authors of AT have advanced the argument that Shannon entropy or LZ compression can only be applied to bits, but this is also mistaken; it can be applied to any object and on any basic units, from vertices in a network to peaks from a mass spectral file, just as they apply their own Ai metric to computer files and computable data representations.
All digital algorithms—including those underpinning AT—operate on representations. Whether it is a mass spectrometry file, a graph of molecular connectivity, or a string representation of a chemical compound, the data is symbolic and abstracted from the physical world. No computer ingests a molecule directly. Not AT, not any other method.
The difference, then, is not in whether abstraction is used, but in how effectively the abstraction captures structure. And here, information theory wins. Compression methods like Shannon entropy or Lempel-Ziv (LZ) not only operate on the same data that AT uses, they do so more efficiently and with mathematically grounded definitions of complexity and redundancy. What AT introduces are additional, ad hoc heuristic layers that obscure rather than clarify the structure being measured.
To make the point more concrete, imagine that a repeating subunit in a molecule, say a benzene ring, is encoded, in a simplified representation, as the binary string 1001. Now consider a polymer such as poly(p-phenylene), composed of multiple such units in a linear chain. Its structure might be symbolically represented as 1001100110011001. A compression algorithm like LZ will quickly identify the repeated pattern and reduce the representation by storing the motif once and indexing its repetitions. From the compressed structure, the original repeating unit 1001 can be fully reconstructed. Rather than broken down, the underlying molecular structure is recovered from the symbolic pattern, even if it has been broken into smaller parts or encoded in a non-chemical alphabet. However, no description of a statistical algorithm includes as a necessary condition that dictionary units or tokens have to be binary code. It can be any alphabet, including the alphabet of molecules that can be forced to not be broken down any further.
The essential insight here is that structure and repetition transcend the encoding format. This is exactly what Shannon formalised in his theory of communication: that the information content of a message is determined by the statistical properties of its components, not by the medium or alphabet used. As long as the representation is lossless and computable, the patterns it contains are physically meaningful and recoverable independent of vocabulary. Turing’s universality and algorithmic complexity results extend this principle, showing that any computable transformation, including molecular assembly, can be simulated symbolically with no loss of information.
Yet, one of the inputs used in AT is mass spectrometry data—one of many possible encodings of a molecule’s structure. In whatever format it is, it is not the molecule itself, but always a digital representation of it. In this respect, AT does not operate closer to the “physical” object than any compression algorithm.
Ultimately, AT’s inability to demonstrate superiority over existing measures, combined with its refusal to engage in rigorous comparisons, undermines its validity. Its claims to measure complexity in a novel way remain unsupported, and its insistence on dismissing well-established metrics such as Shannon entropy and LZ compression on the basis that AT explains selection and evolution, despite a lack of any evidence, cannot withstand basic scrutiny.
3.7 Determinism and Uncertainty
It may seem paradoxical to think that a theory of Turing machines such as algorithmic probability can act or represent a framework for reasoning under uncertainty, when built entirely on deterministic Turing machines or a universal Turing machine (UTM). Yet, this is precisely what gives it its power. A universal Turing machine, although fully deterministic, can generate all possible computable explanations. The probability assigned to each is then determined by both continuous matching with an unfolding observation and, following algorithmic probability [10], the cumulative weight of the number of short programs that reproduce it, the ultimate accepted Bayesian mathematical theory of induction.
The key insight is that uncertainty is not ontological, but epistemic, just as it is not in the real world (e.g. tossing a coin is fully deterministic, we just model it as a random variable out of our own ignorance and limitations). A Turing machine (TM) itself may not be random, but our ignorance of which specific program may have produced the data introduces the element of uncertainty that is quantified and a confidence value assigned. A UTM that is capable of matching or assigning a new computable model (a TM) to every new piece of information produced, thereby updating its internal belief, functions as a model for scientific induction [10]. Rather than modelling stochastic noise or correlations, algorithmic probability ranks explanations in terms of their algorithmic plausibility. It is not about statistical random variables or curve-fitting; it is about causal adequacy and descriptive simplicity. This makes algorithmic probability the optimal solution to the problem of induction, as it formalises the most powerful form of abduction: reasoning to the best explanation. The framework is entirely grounded in the universality of Turing machines, which guarantees that no computable hypothesis space is excluded.
3.8 On the layered accumulation of conceptual and technical flaws and errors
The missteps made by the authors of AT stack on one another in a way that is difficult to peel back. The argument advanced to defend the metric not only collapses under its own weight—since, if valid, it would apply equally against their own measure—but also inverts the very role of universality. Universality, whether in computation, construction, or logic, was introduced precisely to dissolve dependence on any particular machine or substrate. To this is added a contrast with Kolmogorov complexity built on a false problem that von Neumann never engaged—the supposed requirement of physical construction—when in fact his universal constructor was only ever a formal existence proof. The outcome is not only technically incorrect and methodologically limited, but also philosophically confused: abstractions are invoked in ways that reverse their original intent.
When framed against the universality of computation and the optimality of algorithmic probability, Assembly Theory is not merely redundant; it is an unnecessary detour that mistakes statistical regularity for causal explanation.